OpenStudy (anonymous):

Given the linear equation y=-2/7x+3, find the y-coordinates of the points (-7, ), (0, ), and (7, ). Please show all of your work

OpenStudy (aroub):

You have an ordered pair (x,y) you're given x and you have to find y, how? all you gotta do is substitute the x in the given linear equation and you'll find y :) I'm going to the first one to show you how.. The given linear equation: y=-2/7x+3 and an ordered pair (-7,y) substitute 7 by x: y=-2/7(-7)+3 y=2+3 y=5 We found y! The first ordered pair is (-7,5) Can you do the rest on your own? :)

OpenStudy (anonymous):

how did you get the 2+3 to =5

OpenStudy (aroub):

Come on!

OpenStudy (anonymous):

is it cuz the -7+7=0 so that leaves the 2+3

OpenStudy (aroub):

No,it's actually like this: 

OpenStudy (anonymous):

thats what i put